3.3) Orders for jeans from a particular manufacturer for this years Christmas shopping season must be placed in June. The cost per unit for a particular dress is $20 while the selling price is $60. Projected demand is 50, 60, or 70 units. There is a 40 percent chance that demand will be 50 units, a 50 percent chance that demand will be 60 units, and a 10 percent chance that demand will be 70 units. The company believes that any leftover jeans will be donated. How many units should be ordered in June? (Calculations and formulas for all data in the Decision Table must be provided for credit. Answers not in Excel QM and without calculations and formulas for the data will not receive credit.)
Answer #1

Calculation of Jeans ordered in june

Probability of demand is 50 units = 0.4

probability of demand is 60 units = 0.5

probability of demand is 70 units = 0.1

projected demand is = 50*0.4+60*0.5+70*0.1

= 20+30+7

= 57

Decision table

Particulers 50 units 60 units 70 units
Selling price 60 60 60
cost 20 20 20
profit per unit 40 40 40
demand 57 57 57
Remining -7 3 13

if 50 units are ordered : projected demand is 57 units so we are missed the profit of 7*40(60-20)= $ 280 units

if 60 units are ordered : projected demand is 57 units so 3 are leftout and donated so extra  cost is = $ 20*3= $60

if 70 units are ordered : projected demand is 57 units so 13 units are leftout and donated so extra cost is = $ 20*13=$ 260

based on above 60 units are better choice